the sum of four consecutive integers is 40 less then five times the third integer.
Answer by CubeyThePenguin(3113)

consecutive integers: x, (x+1), (x+2), (x+3)

x + (x+1) + (x+2) + (x+3) = 5(x+2) - 40
4x + 6 = 5x - 30
36 = x

The integers are 36, 37, 38, and 39.